# Quantum Machine Learning: from NISQ to Fault Tolerance

### v1

@@ IV.2.1 Machine learning with classical shadow
In summary, classical shadows serve as an efficient quantum-to-classical converter, enabling the extraction of classical information from quantum states. This information can be used to predict observable expectation values. When considering the Variational Quantum Eigensolver (VQE) in a classical context, it resembles a kernel method. The process involves initiating with parameters that define the circuit, creating a map in a high-dimensional space. These parameters are then used to map to a state generated by the circuit, constructing an objective function with the Hamiltonian to acquire classical information. Moreover, classical shadows have demonstrated significant performance gains for VQE Elben et al. (2022), prompting exploration into their potential implementation for more sophisticated machine learning programs.
@@
The central concept involves using classical shadows as a connection between quantum training data and classical predictions. This approach facilitates the learning of ground state properties for Hamiltonians that were not part of the initial training set. The goal is to establish an efficient representation for future Hamiltonians, enabling the computation of observable expectation values and making predictions. This program has been demonstrated to be effective. In Huang et al. (2022b), the authors established that for a small average case prediction error (averaged over all possible inputs), a scenario common in machine learning, the required amount of training data is not extensive. Instead, it scales quasi-polynomially in the number of inputs, not the number of qubits.
@@
In the study conducted by Huang et al. (2021a), the authors delved deeper into the efficacy of this method. They utilized a quantum device to generate data, employed classical shadows to effectively utilize the quantum data, and then applied classical machine learning algorithms for training. They compared this approach (classical machine learning with quantum data leading to classical shadows) to a hypothetical full-scale quantum machine learning protocol. The latter runs on a quantum computer and has coherent access to training data through QRAM. This is the most powerful program that we can envision in the future of QML. The results indicated that, when focusing on average error, there is no significant separation between quantum and classical training data size in terms of sample complexity. However, for worst-case prediction error, Huang et al. (2021a) presented concrete examples demonstrating an exponential separation in complexity. In the research presented in Huang et al. (2022a), the authors explore analogous approaches to learning. They investigate how classical and quantum agents, given their distinct methods of receiving, processing, and storing information, differ in efficiency when aiming to characterize a physical system or state $\rho$. The study demonstrates that quantum agents can achieve exponential advantages in experiments, particularly in terms of the number of queries required to obtain $\rho$.
IV.3 Quantum machine learning for quantum data and quantum simulators
@@
The most immediate application of quantum machine learning lies in handling quantum data Biamonte et al. (2017), i.e., the states produced by quantum systems and processes. Traditional quantum machine learning algorithms, as discussed earlier, identify patterns in classical data by transforming it into quantum states and manipulating those states using basic quantum linear algebra operations. These same algorithms can be directly employed on quantum states of light and matter to uncover inherent features and patterns. The quantum methods of analysis often prove to be more efficient and insightful than classical approaches when dealing with data obtained from quantum systems.
@@
For instance, in the case of multiple copies of a system represented by an $N\times N$ density matrix, quantum principal component analysis can determine eigenvalues and reveal corresponding eigenvectors in a time complexity of $\mathcal{O}\left(\left(\log_{2}{N}\right)^{2}\right)$. This stands in contrast to the $\mathcal{O}(N^{2})$ measurements required for classical devices to perform tomography on the density matrix and the $\mathcal{O}(N^{2}$) operations needed for classical PCA. This quantum analysis of quantum data can be efficiently conducted on the relatively modest-scale quantum computers expected to be available in the coming years. As mentioned before, Huang et al. (2022a) also demonstrated the quantum advantages of learning a physical state $\rho$ through a quantum agent for example quantum sensor or simulator discussed above.
@@
Quantum simulators serve as powerful tools for analyzing quantum dynamics. These are essentially “quantum analog computers” capable of emulating the dynamics of specific quantum systems. Quantum simulators can be specialized devices designed for simulating particular classes of quantum systems or general-purpose quantum computers. The approach involves connecting a trusted quantum simulator to an unknown system, adjusting the simulator’s model to counteract the unknown dynamics, and efficiently learning the dynamics of the unknown system using approximate Bayesian inference Biamonte et al. (2017); Granade et al. (2012); Wiebe et al. (2014, 2015). This significantly reduces the number of measurements required for simulation, offering an exponential improvement.
@@
Notably, the universal quantum emulator algorithm Marvian and Lloyd (2016) enables the reconstruction of quantum dynamics, and the quantum Boltzmann training algorithm Biamonte et al. (2017) allows for the reconstruction of states with a time complexity logarithmic in the dimension of the Hilbert space. This is significantly faster than reconstructing dynamics through classical tomography.
@@
While using a quantum computer to characterize a quantum system Wiebe et al. (2014, 2015) or input states for quantum PCA poses technical challenges, as it doesn’t require QRAM, it holds promise for near-term applications of quantum machine learning Wiebe et al. (2014, 2015); Biamonte et al. (2017), offering the potential for significant speedups in device characterization.
V Conclusion and discussion
@@
This review highlights that both current NISQ-era quantum devices and future fully fault-tolerant quantum computers (FTQC) hold significant promise for applications in machine learning and data analysis Bharti et al. (2021); Biamonte et al. (2017); Liu et al. (2022c, 2023). Over the past decade, quantum computing has witnessed notable advancements in applications, experimental validations, and theoretical findings. The quantum computation field, especially in NISQ applications, has seen an almost exponential growth in the number of research papers. Various factors contribute to this trend, including substantial enhancements in quantum hardware Bharti et al. (2021); Huang et al. (2022a). Given that quantum computing is a relatively young scientific discipline, there is ample opportunity for groundbreaking research and discoveries. The presence of theoretical, practical, and experimental challenges, many of which are addressed in this review, further underscores the motivation for an open-source approach in the field.
@@
As mentioned in Bharti et al. (2021), we expect experimental pursuit in the NISQ era would focus on the design of quantum hardware with a larger number of qubits, and gates with lower error rates capable of executing deeper circuits. Along the way, one of the goals is to demonstrate quantum advantage for practical use cases. If the NISQ paradigm is not powerful enough to exhibit any quantum advantage, theoretical pursuits would be required to understand its limitations. The prime direction of the NISQ and near-term era is to engineer the best possible solution with the limited quantum resources available. The tools and techniques invented during this period could be valuable in the fault- tolerant era as well. To conduct a successful demonstration of quantum advantage, the right blend of the two crucial components: hardware and algorithms designs, is required Bharti et al. (2021). First, hardware development is the key. The design of quantum computers with more qubits, lesser error rates, longer coherence times, and more connectivity between the qubits will be one of the top priorities in the NISQ era. Intensive research in new qubits developments, quantum optimal control and material discovery will be indispensable for both universal programmable quantum computers or special purpose ones. Secondly, to harness the potential of noisy but powerful quantum devices, we expect breakthroughs on the algorithm frontier. Algorithms with realistic assumptions, as the ones mentioned in Bharti et al. (2021), regarding device capabilities will be favored. To lessen the effect of noise, progress towards the design of error suppression, mitigation and correction methods is expected.
@@
While QML has been proposed as a potential avenue for achieving scientific value in the near term using NISQ devices, questions arise regarding its applicability in the future Cerezo et al. (2022a). Researchers envision two distinct post-NISQ eras. The first, termed “partial error corrected,” assumes that quantum computers will possess a sufficient number of physical qubits (a few hundred) and low error rates, allowing for a small number of fully error-corrected logical qubits. In this era, users can allocate qubits between error-corrected and non-error-corrected subsets. The subsequent era, labeled the “fault-tolerant era,” will emerge when quantum hardware features a large number of error-corrected qubits. In this era, algorithms like those discussed in section III.3.1 could notably enhance the scalability and sustainability of classical large-scale machine learning models. Notably, works such as Liu et al. (2021) offer solid theoretical assurances and intersections with cutting-edge classical machine learning research. This approach diverges from the variational quantum algorithms mindset, exemplified by Liu et al. (2022c), by aiming to enhance classical machine learning through a key quantum step that acts as a bottleneck for classical training. Developments in shadow tomography and related fields Aaronson (2018); Elben et al. (2022); Huang et al. (2021a) demonstrate the potential for systematically analyzing QML and gaining a deeper understanding of how quantum speedup is achieved. As demonstrated experimentally in Huang et al. (2022a), with quantum technology such as quantum sensor, quantum memory, and quantum computer, our ability to learn about the physical world might be exponentially improved. Hence, there is confidence that in the future, researchers will have a much more robust understanding of how quantum computing achieves speedup and can develop more powerful applications for machine learning problems.

### v2

@@ 4.2.1 Machine learning with classical shadow
In summary, classical shadows serve as an efficient quantum-to-classical converter, enabling the extraction of classical information from quantum states. This information can be used to predict observable expectation values. When considering the Variational Quantum Eigensolver (VQE) in a classical context, it resembles a kernel method. The process involves initiating with parameters that define the circuit, creating a map in a high-dimensional space. These parameters are then used to map to a state generated by the circuit, constructing an objective function with the Hamiltonian to acquire classical information. Moreover, classical shadows have demonstrated significant performance gains for VQE [41], prompting exploration into their potential implementation for more sophisticated machine learning programs.
@@
The central concept involves using classical shadows as a connection between quantum training data and classical predictions. This approach facilitates the learning of ground state properties for Hamiltonians that were not part of the initial training set. The goal is to establish an efficient representation for future Hamiltonians, enabling the computation of observable expectation values and making predictions. This program has been demonstrated to be effective. In [68], the authors established that for a small average case prediction error (averaged over all possible inputs), a scenario common in machine learning, the required amount of training data is not extensive. Instead, it scales quasi-polynomially in the number of inputs, not the number of qubits.
@@
In the study conducted by [66], the authors delved deeper into the efficacy of this method. They utilized a quantum device to generate data, employed classical shadows to effectively utilize the quantum data, and then applied classical machine learning algorithms for training. They compared this approach (classical machine learning with quantum data leading to classical shadows) to a hypothetical full-scale quantum machine learning protocol. The latter runs on a quantum computer and has coherent access to training data through QRAM. This is the most powerful program that we can envision in the future of QML. The results indicated that, when focusing on average error, there is no significant separation between quantum and classical training data size in terms of sample complexity. However, for worst-case prediction error, [66] presented concrete examples demonstrating an exponential separation in complexity. In the research presented in [67], the authors explore analogous approaches to learning. They investigate how classical and quantum agents, given their distinct methods of receiving, processing, and storing information, differ in efficiency when aiming to characterize a physical system or state $\rho$. The study demonstrates that quantum agents can achieve exponential advantages in experiments, particularly in terms of the number of queries required to obtain $\rho$.
4.3 Quantum machine learning for quantum data and quantum simulators
@@
The most immediate application of quantum machine learning lies in handling quantum data [18], i.e., the states produced by quantum systems and processes. Traditional quantum machine learning algorithms, as discussed earlier, identify patterns in classical data by transforming it into quantum states and manipulating those states using basic quantum linear algebra operations. These same algorithms can be directly employed on quantum states of light and matter to uncover inherent features and patterns. The quantum methods of analysis often prove to be more efficient and insightful than classical approaches when dealing with data obtained from quantum systems.
@@
For instance, in the case of multiple copies of a system represented by an $N\times N$ density matrix, quantum principal component analysis can determine eigenvalues and reveal corresponding eigenvectors in a time complexity of $\mathcal{O}\left(\left(\log_{2}{N}\right)^{2}\right)$. This stands in contrast to the $\mathcal{O}(N^{2})$ measurements required for classical devices to perform tomography on the density matrix and the $\mathcal{O}(N^{2}$) operations needed for classical PCA. This quantum analysis of quantum data can be efficiently conducted on the relatively modest-scale quantum computers expected to be available in the coming years. As mentioned before, [67] also demonstrated the quantum advantages of learning a physical state $\rho$ through a quantum agent for example quantum sensor or simulator discussed above.
@@
Quantum simulators serve as powerful tools for analyzing quantum dynamics. These are essentially “quantum analog computers” capable of emulating the dynamics of specific quantum systems. Quantum simulators can be specialized devices designed for simulating particular classes of quantum systems or general-purpose quantum computers. The approach involves connecting a trusted quantum simulator to an unknown system, adjusting the simulator’s model to counteract the unknown dynamics, and efficiently learning the dynamics of the unknown system using approximate Bayesian inference [18, 48, 132, 133]. This significantly reduces the number of measurements required for simulation, offering an exponential improvement.
@@
Notably, the universal quantum emulator algorithm [96] enables the reconstruction of quantum dynamics, and the quantum Boltzmann training algorithm [18] allows for the reconstruction of states with a time complexity logarithmic in the dimension of the Hilbert space. This is significantly faster than reconstructing dynamics through classical tomography.
@@
While using a quantum computer to characterize a quantum system [132, 133] or input states for quantum PCA poses technical challenges, as it doesn’t require QRAM, it holds promise for near-term applications of quantum machine learning [132, 133, 18], offering the potential for significant speedups in device characterization.
5 Conclusion and discussion
@@
This review highlights that both current NISQ-era quantum devices and future fully fault-tolerant quantum computers (FTQC) hold significant promise for applications in machine learning and data analysis [17, 18, 87, 89]. Over the past decade, quantum computing has witnessed notable advancements in applications, experimental validations, and theoretical findings. The quantum computation field, especially in NISQ applications, has seen an almost exponential growth in the number of research papers. Various factors contribute to this trend, including substantial enhancements in quantum hardware [17, 67]. Given that quantum computing is a relatively young scientific discipline, there is ample opportunity for groundbreaking research and discoveries. The presence of theoretical, practical, and experimental challenges, many of which are addressed in this review, further underscores the motivation for an open-source approach in the field.
@@
As mentioned in [17], we expect experimental pursuit in the NISQ era would focus on the design of quantum hardware with a larger number of qubits, and gates with lower error rates capable of executing deeper circuits. Along the way, one of the goals is to demonstrate quantum advantage for practical use cases. If the NISQ paradigm is not powerful enough to exhibit any quantum advantage, theoretical pursuits would be required to understand its limitations. The prime direction of the NISQ and near-term era is to engineer the best possible solution with the limited quantum resources available. The tools and techniques invented during this period could be valuable in the fault- tolerant era as well. To conduct a successful demonstration of quantum advantage, the right blend of the two crucial components: hardware and algorithms designs, is required [17]. First, hardware development is the key. The design of quantum computers with more qubits, lesser error rates, longer coherence times, and more connectivity between the qubits will be one of the top priorities in the NISQ era. Intensive research in new qubits developments, quantum optimal control and material discovery will be indispensable for both universal programmable quantum computers or special purpose ones. Secondly, to harness the potential of noisy but powerful quantum devices, we expect breakthroughs on the algorithm frontier. Algorithms with realistic assumptions, as the ones mentioned in [17], regarding device capabilities will be favored. To lessen the effect of noise, progress towards the design of error suppression, mitigation and correction methods is expected.
@@
While QML has been proposed as a potential avenue for achieving scientific value in the near term using NISQ devices, questions arise regarding its applicability in the future [27]. Researchers envision two distinct post-NISQ eras. The first, termed “partial error corrected,” assumes that quantum computers will possess a sufficient number of physical qubits (a few hundred) and low error rates, allowing for a small number of fully error-corrected logical qubits. In this era, users can allocate qubits between error-corrected and non-error-corrected subsets. The subsequent era, labeled the “fault-tolerant era,” will emerge when quantum hardware features a large number of error-corrected qubits. In this era, algorithms like those discussed in section 3.3.1 could notably enhance the scalability and sustainability of classical large-scale machine learning models. Notably, works such as [90] offer solid theoretical assurances and intersections with cutting-edge classical machine learning research. This approach diverges from the variational quantum algorithms mindset, exemplified by [87], by aiming to enhance classical machine learning through a key quantum step that acts as a bottleneck for classical training. Developments in shadow tomography and related fields [2, 41, 66] demonstrate the potential for systematically analyzing QML and gaining a deeper understanding of how quantum speedup is achieved. As demonstrated experimentally in [67], with quantum technology such as quantum sensor, quantum memory, and quantum computer, our ability to learn about the physical world might be exponentially improved. Hence, there is confidence that in the future, researchers will have a much more robust understanding of how quantum computing achieves speedup and can develop more powerful applications for machine learning problems.